Write and equation of the line that passes through each pair of points.
(-5/4,1) (-1/4,3/4)

Answer by stanbon(75887)

Write and equation of the line that passes through each pair of points.
(-5/4,1) (-1/4,3/4)
----
slope = (3/4 - 1)/(-1/4 -(-5/4)) = (-1/4)/(1) = -1/4
---------------------
Form: y = mx + b
1 = (-1/4)(-5/4) + b
1 = 5/16 + b
b = 11/16
-----------------
Equation:
y = (-1/4)x + (11/16)
============
Cheers,
Stan H.
============